in a poll of women 18 and older, 70% of women polled felt their financial situation was better than their mothers. assume that this percentage is true fo the current populatiom of all women 18 and older. find the probability that the proportion of women in a random sample of 400 who hold this view is less than .73
Answer by stanbon(57967)

in a poll of women 18 and older, 70% of women polled felt their financial situation was better than their mothers.
-------------
assume that this percentage is true for the current populatiom of all women 18 and older.
-------------
find the probability that the proportion of women in a random sample of 400 who hold this view is less than .73
-------------------------------
z(0.73)=(0.73-0.70)/sqrt(0.7*0.3/400) = 1.30931
------------------------
P(p-hat < 1.30931) = P(z < 1.30931) = 0.9048
=================================================
Cheers,
Stan H.